OpenStudy (anonymous):

Solve the following system using the graphing method: y = -x + 2 y = x + 4

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@IMStuck PICS ATTATCHED

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (imstuck):

I'm here. sorry it took so long.

OpenStudy (anonymous):

its ok

OpenStudy (imstuck):

the solution to this system is where on the graph the lines cross each other. So you need to decide which graph depicts your given lines. Do you know how to determine a graph from an equation in the form y = mx + b?

OpenStudy (anonymous):

not really I have pics attathed by the way

OpenStudy (imstuck):

it's the last graph on pic_21.jpg

OpenStudy (imstuck):

not the one you just attached. Its the other one, the second graph on that second page.

OpenStudy (anonymous):

ok thank :)

OpenStudy (anonymous):

how about this one

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

is the system y = -4x + 1 and -12x + 3y = 3

OpenStudy (anonymous):

no y-4x+1 -12x+3y-3

OpenStudy (imstuck):

oh that changes everything!

OpenStudy (anonymous):

lol

OpenStudy (imstuck):

There are no equal signs in those!!!???

OpenStudy (anonymous):

none

OpenStudy (imstuck):

give me a sec then, ok?

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

in my experience you have to have a y = something line in order to graph. From what I can tell about those lines is that it is the second choice down on the image you posted.

OpenStudy (anonymous):

b?

OpenStudy (aum):

Without equal signs, there are no equations.

OpenStudy (anonymous):

is there a graphingcalculator I can use

OpenStudy (aum):

It is a small print but I think the equations are: y = 4x + 1 (not y minus 4x but y equals 4x + 1) -12x + 3y = 3 (not minus 3 but equal to 3)

OpenStudy (anonymous):

y = 2x + 1 -4x + 2y = 4 would this be inconsistent?

OpenStudy (aum):

y = 4x + 1 -12x + 3y = 3 Take the second equation and solve for y. -12x + 3y = 3 add 12x 3y = 12x + 3 divide by 3 y = 4x + 1 The second equation is identical to the first equation! Since both equations are the same they are consistent and dependent. In the graph the two equations will show up as one single line with a positive slope of 4.

OpenStudy (anonymous):

ok so both are consistent

OpenStudy (aum):

consistent and dependent

OpenStudy (anonymous):

does it look like this @aum

OpenStudy (aum):

look like what?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

?????

OpenStudy (aum):

No. Both equations are the same: y = 4x + 1. So if you plot them there will be just one line. And slope m = 4. So it should have a positive slope.

OpenStudy (anonymous):

all of my choices have two lines

OpenStudy (aum):



OpenStudy (anonymous):

y = 2x + 1 -4x + 2y = 4

OpenStudy (anonymous):

for these

OpenStudy (anonymous):

so that one is C?

OpenStudy (aum):

Yes.

OpenStudy (aum):

y = 2x + 1 -4x + 2y = 4 Follow the same procedure as before. Take the second equation and solve for y. What do you get?

OpenStudy (anonymous):

those ones actually did have equal signs

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

for this one?y = -x + 2 y = x + 4

OpenStudy (anonymous):

y = 2x + 1 -4x + 2y = 4

OpenStudy (aum):

They are two different lines. Are they parallel?

OpenStudy (anonymous):

this one sorry its hard without the actual graph

OpenStudy (anonymous):

no not parallel

OpenStudy (anonymous):

ok which one are we on

OpenStudy (anonymous):

sorry

OpenStudy (aum):

y = -x + 2 y = x + 4 Are these lines parallel? (that is, are their slopes the same?)

OpenStudy (anonymous):

have you get it..?